solve for X and Y using substitution method
2x=y+1
x+y=-1
thank you

Answer by solve_for_x(190)

Solving the first equation for y gives:

y = 2x - 1

Substituting 2x - 1 in place of y in the second equation then gives:

x + (2x - 1) = -1

x + 2x - 1 = -1

3x - 1 = -1

3x = -1 + 1

3x = 0

x = 0

Then, with x = 0

y = 2x - 1 = 2(0) - 1 = -1

Solution: x = 0, y = -1